Thank you in advance! This is Environmental Economics.2. The number of visitors to the zoo and the entrance fee is shown in the table below: 800 600 400 200 0 0 6 9 12 Calculate the change in consumer surplus if the entrance fee increases from three (3) to nine Euro (9). How does the price increase affect the revenue of the zoo?
Answer #1

When fee was € 3

Consumer Surplus = (1/2)*(12 -3)*(600 -0) = € 2700

When entrance fee was increased to € 9

Consumer Surplus = (1/2)*(12 - 9)*(200 -0) = € 300

Change in consumer surplus = CS€9 - CS€3 = 300-2700

Change in consumer surplus = -€ 2400

The negative sign means the Consumer Surplus decreased by € 2400.

REVENUE
Entrance Fee #Visitor's Revenue
0 800 0*800 = € 0
3 600 3*600 = € 1800
6 400 6*400 = € 2400
9 200 9*200 = € 1800
12 0 12*0 = € 0

No change in the revenue.

When entrance fee was €3 then revenue = € 1800

When fee was increased to €9 then revenue = € 1800

Thus, revenue is same in both the cases.

Please contact if having any query thank you.